### v1

# Difference of surface effects in the inclusive and reactions

Abstract
Background: Previous studies have revealed surface effects in inclusive and reactions. These findings have contributed significantly to the improvement of nuclear data evaluation. However, surface effect in inclusive reaction has hardly been investigated.
Purpose: The purpose of this study is to investigate the difference of the surface effects in the and reactions. We also aim to clarify the origin of the difference.
Methods: The energy spectra and their radial distributions for the and reactions are calculated by the one-step semiclassical distorted wave model.
Results: In the region with small energy transfer, the radial distribution of the energy spectra for the reaction is found to be shifted toward the outer region of the nucleus more than that of the reaction. Furthermore, the calculations ignoring the nuclear absorption show that the radial distributions in the two reactions are nearly identical.
Conclusion: The surface effect in the reaction is more pronounced than in the reaction. A significant cause of the difference is that deuteron undergoes stronger absorption by the nucleus than proton.
I INTRODUCTION
Phenomenological models have played an important role in the nuclear data evaluation for nucleon-induced reactions [1, 2, 3, 4, 5]. For the calculation of pre-equilibrium processes, which are prominent at incident energies above about 10 MeV, the exciton model [6] has been widely used. This phenomenological model has achieved great success in combination with global parameterization based on analyses of a large number of experimental data for reactions [7, 8]. It has been argued in these analyses that the inclusion of a surface localization effect is important in reproducing the shape of experimental energy spectrum [7, 9]. This effect is taken into account by introducing corrections to the density of states in the exciton model. Note that the importance of the surface effect is supported not only by the comparison with experimental data but also by the microscopic theoretical analysis [10] using the semiclassical distorted wave (SCDW) model, which has no free adjustable parameter [11, 12, 13, 14, 15].
Evaluation activities of deuteron nuclear data have gradually begun [1, 16, 4], mainly motivated by the development of deuteron accelerator-based intensive neutron sources [17, 18]. In the evaluation of deuteron nuclear data, phenomenological models [19, 20] using state densities similar to those in the exciton model are often employed to calculate the components of inelastic scattering to continuum states. The evaluation of the inelastic scattering is important since it affects the transport of deuterons in a material. Therefore, as is the case with the exciton model, whether the surface effect should be taken into account in the state densities used in the above inelastic scattering model, and if so, to what extent, is an important factor in the deuteron nuclear data evaluation. However, since experimental data for reactions are very limited, there is concern that making this judgment based on a comparison with specific experimental data may give unphysical results under other conditions. A deep and comprehensive understanding for surface effects in reactions is required.
Under these circumstances, in Ref. [21], the one-step SCDW model [11] has been successfully applied to the description of inclusive reactions. The SCDW model has the property which describes the inclusive cross section as an incoherent integral of the contributions at individual collision points [22, 10]. Therefore, the model is expected to be useful for investigating the peripherality of the reaction. Findings from the microscopic reaction model analysis will provide guidance and support for the consideration of the surface effect in the phenomenological deuteron inelastic scattering models.
The purpose of this study is to investigate the surface effect in the reaction and to clarify the difference from that in the reaction. We also aim to clarify the cause of the difference. The above analyses are performed utilizing the above-mentioned property of the SCDW model. Exploring the difference in surface effects between the two reactions are of interest from the view point not only of nuclear data evaluation but also of fundamental nuclear physics. Furthermore, we apply the findings from the SCDW analysis to the implementation of the surface effect into the phenomenological deuteron inelastic scattering model. This is intended to show how much microscopic reaction theory contributes to improving phenomenological models.
The construction of this paper is as follows. In Sec. II we briefly describe the one-step SCDW model for the inclusive and reactions. We also explain here the phenomenological inelastic scattering model. In Sec. III we compare the calculated energy spectrum of the inclusive and reactions with experimental data and clarify the difference of the surface effects between the two reactions. Applications for improving the phenomenological model also be conducted here. Finally, a summary is given in Sec. IV.
II Methods
II.1 SCDW model
We briefly describe the inclusive and reactions with the one-step SCDW model. The double differential cross section for the energy and the solid angle of the emitted particle is expressed with
where and are the mass numbers of the particle and the target nucleus, respectively. is the asymptotic momentum of the incident (emitted) particle, is the coordinate of the collision point with respect to the center of the target nucleus. The distorted waves for in the initial and final states are denoted by and , respectively. is the nuclear density at . The averaged double differential cross section of the elementary process at is given by
where the subscript denotes the label of proton and neutron in the target. is the momentum of the nucleon in the target in the initial state. The local Fermi momenta of nucleon and are denoted by and , respectively. is the free scattering cross section determined by the local scattering angle and the local scattering energy between and the nucleon in the target. is the incident energy of and is the kinetic energy of the nucleon in the target nucleus in the initial (final) state.
The energy spectra of the and reactions are obtained by integrating Eq. (1) over :
By applying the local Fermi gas model (LFG) to the initial and final nuclear single-particle states, is given by [13]
where and are atomic and neutron numbers of the target nucleus, respectively. LFG can takes into account the spread of nucleon density on the nuclear surface because the local Fermi momentum depends on . To clarify the surface effect, we also consider the Fermi gas model (FG), which give a uniform Fermi momentum distribution. When FG is used instead of LFG, and become step functions with respect to in Eq. (2). On the other hand, the Fermi momentum in LFG decreases smoothly as increases, as shown later in Fig. 1.
We use the proton- and deuteron-nucleus global optical potential by Koning-Delaroche [23] and An-Cai [24] for describing projectile-nucleus scattering, respectively. The nuclear density is assumed to be the Woods-Saxon form, where the radial parameter is defined as , with = 1.15 fm, and the diffuseness parameter is set to fm as in Ref. [21]. For the differential cross sections of - scattering used in the calculation, we utilize the numerical table from Ref. [25], which was made by fitting experimental data of - elastic scattering. In the calculations of the process, the free - scattering cross sections are calculated by using the nucleon-nucleon matrix provided by Franey and Love [26, 27].
II.2 Phenomenological model by Kalbach
As a phenomenological model for continuum deuteron inelastic scattering, we adopt the model proposed by Kalbach [28]. Although this model was originally developed for -particle-induced reactions, our previous study has shown that the model is applicable to deuteron-induced ones [19].
According to Ref. [28], the energy spectrum of the deuteron inelastic scattering with a nucleon particle-hole pair creation is calculated by:
where is a normalization constant determined from fitting experimental energy spectrum, is the deuteron total reaction cross section at scattering energy , and and are the incident and emitted deuteron energies, respectively. is obtained from the optical model calculation with the global potential by An-Cai [24]. is the spin of the emitted deuteron, respectively. denotes the mass number of the target, and represents the excitation energy of residual nucleus. The final state density is expressed as follows:
where , , and and are the numbers of neutrons and protons in the target. The setting of values for and follows the original paper by Kalbach [28]. The squaring of and corresponds to considering the - states of nucleons. The finite well function with the well depth is defined for the - states as follows:
where is the unit step function.
A simple method to include surface effects in the exciton model was proposed in Ref. [7], and has been successfully applied to the analysis of reactions [9, 8]. In the method, the surface effects are taken into account by setting the small value of for the =1 state. In this work we apply this method to Eq. (6). In other words, we set in Eq. (8) smaller than the typical value of 38 MeV associated to the Fermi energy. The details of setting the value of are described later in Sec. III.3.
III RESULTS AND DISCUSSION
III.1 Effect of momentum distribution of nucleons on energy spectrum
We compare the energy spectra of inclusive and reactions calculated with SCDW using LFG and FG, and experimental data. Figure 1 shows the radial distributions of the Fermi momenta of calculated with LFG (solid line) and FG (dashed line).
In FG, no reaction is allowed in fm because is zero in the region. On the other hand, in LFG, reactions are allowed beyond the boundary because LFG incorporates the effect of diffuseness of the nucleus. Therefore, the difference of the SCDW calculations with LFG and FG can be tied to surface effects.
Figure 2 shows the energy spectra of the and reactions for several incident energies and targets. The experimental data are taken from Refs. [29, 30]. In order to make the velocity of projectile same, the incident energies per nucleon in Fig. 2(e) and (f) are the same as those in (b) and (c), respectively. Note that these results are calculated with SCDW which has no free adjustable parameter.
For the reactions, the energy spectra calculated with SCDW using LFG reproduces the experimental data well except for the region with large energy transfer as shown in Fig 2(a), (b), and (c). For the reactions, the energy spectra calculated with LFG reasonably reproduce the experimental data in the region with MeV, as shown in Fig. 2(d). In both reactions, the calculated energy spectrum undershoot the experimental data in the region with low emission energies. This is because the contributions of particle emission from the multi-step direct process and the compound process are dominant in that region [22, 19]. For this reason, in what follows, we will discuss the energy spectra in the region with MeV, where the one-step process is dominant and the elastic scattering events are not included.
By comparing the energy spectra calculated with LFG and FG, we can see that consideration of the surface effects is necessary to reproduce experimental data in both of the and reactions. The surface effects do not depend so much on the incident energy and target nucleus in both reactions. On the other hand, the reactions have stronger dependence than that of the reactions.
III.2 Difference of radial distributions of and reactions
To analyze the difference of the surface effects between the and reactions in more detail, we discuss the radial distribution of the energy spectra.
The radial distribution of the energy spectra is defined by
where is the solid angle of . Note that satisfies .
Figure 3 shows the in the calculation with LFG for (a) the reaction at 62 MeV and (b) the reaction at 124 MeV (62 MeV per nucleon). The solid, dashed, and dotted lines represent with MeV, respectively. The vertical dash-dotted lines indicate fm, where the Fermi momentum of is zero in FG.
As illustrated in Fig. 3, for a given , the peak position is almost the same for the and reactions and the peak positions shift outwards as decreases. In contrast, the dependence of the peak heights of the two reactions is opposite. The peak height of the reaction becomes smaller as decreases, while that of the reaction becomes larger as decreases.
To clearly show how this significant difference affects the surface effects of these reactions, we present in Fig. 4 the integrated values of in the range of 5 15 MeV. The solid and dashed lines represent the calculated values of the and reactions, respectively. Each of them is normalized so that the integrated value over is unity. It is seen that the radial peak position is about 4.3 fm for the reaction. On the other hand, for the reaction, the peak position is about 6.0 fm and the large proportion of reaction occurs in the outer region of the nucleus. This clearly shows that the surface effect is more significant for the reaction.
Next, we discuss the cause of the difference in the surface effects between the two reactions. Figure 5 is the same as Fig. 4 but for the calculations ignoring the absorption of the protons and deuterons by the nucleus; the imaginary part of the optical potential is set to zero in the calculations. When we ignore the absorption, the radial distributions of the two reactions are almost identical. This means that the stronger surface effect in the reaction than in the reaction is attributed to the strong absorption of deuteron by the nucleus. We can also see that, in Fig 3, the strong absorption of the deuterons suppresses the reaction in the inner region of the nucleus in the than in the reaction.
III.3 Application to phenomenological model
Next, we apply the findings obtained from the previous sections to the implementation of the surface effect into the phenomenological model. Figure 6 illustrates the comparison of the experimental data and the energy spectra calculated with the Kalbach model presented in Sec II.2. The experimental data are obtained from Ref. [30].
Three lines are shown in the figure. The solid line is the result of the calculation in which the well depth is optimized to reproduce the experimental spectrum. The optimised value of is 10 MeV. The normalization factor is adjusted so that the peak of the spectrum matches the experimental data. This adjustment holds true also for the two calculations below. The dashed line is the result with the parameterization optimized for proton-induced reactions [8]. In Ref. [8], the analyses using the exciton model were performed for reactions up to 200 MeV for various targets, and the different parameterizations for neutron- and proton-induced reactions were given based on the results. The value of in the case of proton-induced reactions is given as follows:
where and are the incident energy and the target mass, respectively. In the case shown in Fig. 6, Eq. (10) gives about 25 MeV as the value of . The dotted line is the calculation with a typical well depth of 38 MeV. Using this value of corresponds to not considering the surface effects.
As shown in the figure, the calculation not considering the surface effects does not reproduce the shape of the experimental spectrum. Note that the components below 30 MeV are mainly due to the pre-equilibrium process for other states than - and the compound nucleus process. The large component slightly below 80 MeV is attributed to inelastic scattering to low-lying discrete levels and elastic scattering. These components are not considered in the Kalbach model and are thus outside the scope of the present discussion.
In terms of the calculations considering the surface effects, the use of the optimized value for proton-induced reactions improves the agreement with experimental data, but underestimation in the high-energy region is still seen. On the other hand, when the value of is decreased as our SCDW analysis suggests and the value is set to 10 MeV, the calculation results reproduce the experimental energy spectrum well over almost the entire energy range to be considered in the Kalbach model. These results indicate that parameter optimization of the phenomenological Kalbach model has been justified also by theoretical analysis with the microscopic SCDW model.
IV SUMMARY
We have clarified the difference of the surface effects between the inclusive and reactions and the cause of the difference. The energy spectra calculated by SCDW with LFG and FG were compared to the experimental data of the and reactions on several combinations of targets and incident energies. By comparing the energy spectra with LFG and FG, we have found that the surface effect is necessary to reproduce the experimental data for both and reactions.
We have investigated in more details the difference of the surface effects in the two reactions by comparing the radial distributions of the energy spectra. The radial peak positions for each energy transfer shift to the outer region of the nucleus as decrease and the tendencies were almost the same for the two reactions. However, the opposite trend was observed between the two reactions in terms of the peak height. As a result, it has been found that the surface effect becomes larger for the reaction than the reaction. Moreover, we compared the radial distributions ignoring the absorption by the nucleus. The results show almost identical radial distributions and it has been found that the cause of the stronger surface effect in the reaction is the stronger absorption of deuteron than proton.
We have applied the above findings on surface effects in reaction to the improvement of the phenomenological model by Kalbach to calculate continuum deuteron inelastic scattering. We have newly introduced a surface effect into the model. For the finite well depth , which is the adjustable parameter associated with the surface effect, the optimized value for the proton-induced reactions was insufficient to reproduce the experimental data of the reaction. When the value of was set much smaller, as suggested by the analysis with SCDW, the calculated values reproduced the experimental data well over a wide emission energy range. This result has indicated that parameter optimization of a phenomenological model has been justified by a microscopic reaction model.
As one of the future works, it will be necessary to extend the present one-step SCDW model to describe the multi-step processes. With this extension, the SCDW model could guide the improvement of phenomenological models not only with respect to surface effects, which are prominent in the small region, but also with respect to wider energy spectra and angular distributions of outgoing deuterons. On another front, it is of interest to perform a similar analysis for other particles. -particle is a good candidate since it is well known to undergo strong absorption and is of large importance in the application fields.
Acknowledgements.
The authors thank Y. Chazono for providing us with the table of the - scattering cross section, and thank S. Kawase for fruitful discussions. This work has been supported in part by Grants-in-Aid of the Japan Society for the Promotion of Science (Grants No. JP20K14475, No. JP21H00125, and No. JP21H04975). The computation was carried out with the computer facilities at the Research Center for Nuclear Physics, Osaka University.References
- Iwamoto et al. [2023] O. Iwamoto, N. Iwamoto, S. Kunieda, F. Minato, S. Nakayama, Y. Abe, K. Tsubakihara, S. Okumura, C. Ishizuka, T. Yoshida, et al., J. Nucl. Sci. Technol. 60, 1 (2023).
- Brown et al. [2018] D. A. Brown, M. Chadwick, R. Capote, A. Kahler, A. Trkov, M. Herman, A. Sonzogni, Y. Danon, A. Carlson, M. Dunn, et al., Nucl. Data Sheets 148, 1 (2018).
- Plompen et al. [2020] A. J. Plompen, O. Cabellos, C. De Saint Jean, M. Fleming, A. Algora, M. Angelone, P. Archier, E. Bauge, O. Bersillon, A. Blokhin, et al., Eur. Phys. J. A 56, 1 (2020).
- Koning et al. [2019] A. Koning, D. Rochman, J.-C. Sublet, N. Dzysiuk, M. Fleming, and S. Van der Marck, Nucl. Data Sheets 155, 1 (2019).
- Chadwick et al. [1999] M. Chadwick, P. Young, S. Chiba, S. Frankle, G. Hale, H. Hughes, A. Koning, R. Little, R. MacFarlane, R. Prael, et al., Nucl. Sci. Eng. 131, 293 (1999).
- Kalbach [1986] C. Kalbach, Phys. Rev. C 33, 818 (1986).
- Kalbach [1985] C. Kalbach, Phys. Rev. C 32, 1157 (1985).
- Koning and Duijvestijn [2004] A. Koning and M. Duijvestijn, Nucl. Phys. A 744, 15 (2004).
- Kalbach [2000] C. Kalbach, Phys. Rev. C 62, 044608 (2000).
- Watanabe et al. [1999] Y. Watanabe, R. Kuwata, S. Weili, M. Higashi, H. Shinohara, M. Kohno, K. Ogata, and M. Kawai, Phys. Rev. C 59, 2136 (1999).
- Luo and Kawai [1991] Y. L. Luo and M. Kawai, Phys. Rev. C 43, 2367 (1991).
- Kawai and Weidenmüller [1992] M. Kawai and H. A. Weidenmüller, Phys. Rev. C 45, 1856 (1992).
- Ogata et al. [1999] K. Ogata, M. Kawai, Y. Watanabe, S. Weili, and M. Kohno, Phys. Rev. C 60, 054605 (1999).
- Weili et al. [1999] S. Weili, Y. Watanabe, M. Kohno, K. Ogata, and M. Kawai, Phys. Rev. C 60, 064605 (1999).
- Ogata et al. [2002] K. Ogata, Y. Watanabe, S. Weili, M. Kohno, and M. Kawai, Nucl. Phys. A 703, 152 (2002).
- Nakayama et al. [2021] S. Nakayama, O. Iwamoto, Y. Watanabe, and K. Ogata, J. Nucl. Sci. Technol. 58, 805 (2021).
- Moeslang et al. [2006] A. Moeslang, V. Heinzel, H. Matsui, and M. Sugimoto, Fusion Eng. Des. 81, 863 (2006).
- Ledoux et al. [2014] X. Ledoux, M. Aïche, M. Avrigeanu, V. Avrigeanu, L. Audouin, E. Balanzat, B. Ban-détat, G. Ban, G. Barreau, E. Bauge, et al., Nucl. Data Sheets 119, 353 (2014).
- Nakayama et al. [2022] S. Nakayama, O. Iwamoto, Y. Watanabe, and K. Ogata, Few-Body Syst. 63, 4 (2022).
- Kalbach [2005] C. Kalbach, Phys. Rev. C 71, 034606 (2005).
- Nakada et al. [2023] H. Nakada, K. Yoshida, and K. Ogata, Phys. Rev. C 108, 034603 (2023).
- Cowley et al. [2000] A. A. Cowley, G. F. Steyn, Y. Watanabe, T. Noro, K. Tamura, M. Kawabata, K. Hatanaka, H. Sakaguchi, H. Takeda, and M. Itoh, Phys. Rev. C 62, 064604 (2000).
- Koning and Delaroche [2003] A. Koning and J. Delaroche, Nucl. Phys. A 713, 231 (2003).
- An and Cai [2006] H. An and C. Cai, Phys. Rev. C 73, 054605 (2006).
- Chazono et al. [2022] Y. Chazono, K. Yoshida, and K. Ogata, Phys. Rev. C 106, 064613 (2022).
- Love and Franey [1981] W. G. Love and M. A. Franey, Phys. Rev. C 24, 1073 (1981).
- Franey and Love [1985] M. A. Franey and W. G. Love, Phys. Rev. C 31, 488 (1985).
- Kalbach [1977] C. Kalbach, Z. Phys. A 283, 401 (1977).
- Bertrand and Peelle [1973] F. E. Bertrand and R. W. Peelle, Phys. Rev. C 8, 1045 (1973).
- Wu et al. [1979] J. R. Wu, C. C. Chang, and H. D. Holmgren, Phys. Rev. C 19, 370 (1979).